Let f(x)=16-x² and g(x)=4-x. Find f(x)+g(x).

Answer by Edwin McCravy(20059)



Just add their right sides:

f(x) + g(x) = "right side of f(x)" + "right side of g(x)" 

f(x) + g(x) = (16-x²) + (4-x)

Now simplify:

f(x) + g(x) = 16 - x² + 4 - x

f(x) + g(x) = 20 - x² - x

Arranging in descending order

f(x) + g(x) = -x² - x + 20

Edwin